A person connects four 1/4 Ohm cells in series but one cell has its terminal reversed. the external resistance is 1 Ohm. If Each cell has an emf of 1.5 V the current flowing is -
 
a) 4/3 A
b) 3/4 A
c) 1.5 A
d) zero

Grade:10

4 Answers

siddharth gupta
9 years ago
resistance of all the cells connected in series gets added up with the external resistance to give an effective resistance of 2ohms.The emf of three cells in series with same polarity gets added up to give resultant emf of 4.5V.Now the emf of the cell connected with opposite polarity is subtracted to give the effective emf of 3V.
3V/2ohm=1.5Amp which is the electric current in the circuit.
Ashwin
9 years ago
your answer is correct but 
i want to  know that instead of potential difference you used emf in ohm’s law . shou;ld’nt we have to calculate P.D first in order to calculate current.
Akanksha Das
6 years ago
effective emf is 4.5V/2 ohm=2.25A.Due to opposite polarity 3 cells added i.e.4.5V
If it is same polarity then emf will 6v & for one cell opposite polarity 6V-1.5V=4.5V
Akanksha
Aman Arya
6 years ago
Here four 1/4ohm cells are connected in series with external resistance 1ohmSo effective resistance=1/4+1/4+1/4+1/4+1 =2ohmNow each cell has an e.m.f of 1.5 volt But one cell has its terminal reversed So it exert negative force So total p.d= 1.5+1.5+1.5-1.5=3volt. And then the current flow =V/R =3/2=1.5 AAman Arya